1. Find the positive eigenvalues and the corresponding eigenfunctions of the boundary value problem - +42y = 0; y(0) = 0, y( please show work differential equation
Answer #1

initial Value Problem 16, Given y+4ay = Yo)= 0 , Y(2T)=0 to check that the given Problem NOu bas Positive eigenvelues or noty2) + C26in 4T7 JA C4 cos 4 TI = Prom abeme C6in 4TT o if get tivial 6olution Co=0 we consider c+0, Bin 4T7 4TA TT n=1,2, =1,